OpenStudy (anonymous):

Factor (a^4-b^4) and a^2-2ab+b^2

OpenStudy (cwrw238):

(a^4-b^4) = (a^2 - b^2)(a^2 + b^2) = (a - b)( a + b)(a^2 + b^2)

OpenStudy (anonymous):

Thank you!

OpenStudy (cwrw238):

yw